OpenStudy (anonymous):

limit |x-1|/(x^2+x-2) how to solve by deffinition x->+1 of limits

myininaya (myininaya):

1+?

myininaya (myininaya):

\[x^2+x-2=(x+2)(x-1) ; |x-1|=x-1 \text{ since } x>1\]

OpenStudy (anonymous):

one from above maybe?

OpenStudy (anonymous):

cause limit as 1 does not exist

myininaya (myininaya):

right

myininaya (myininaya):

since right limit does not equal left limit

OpenStudy (anonymous):

The limit doesn't exist at 1 or at -2.... But the limit from the right does exist at 1 as indicated by the above responses.